OpenStudy (anonymous):

Use mathematical induction to prove the statement is true for all positive integers n, or show why it is false. 4 * 6 + 5 * 7 + 6 * 8 + ... + 4n(4n + 2) = 4(4n + 1)(8n + 7) {over} 6

OpenStudy (turingtest):

try this for n=1

OpenStudy (anonymous):

do I put it in the fraction?

OpenStudy (anonymous):

wait, okay I got 24 when I subbed it in and got 24, the same number as when you multiply the first two numbers @TuringTest

OpenStudy (amistre64):

4n(4n + 2) = 4(4n + 1)(8n + 7)/6 ; try n=1, assuming your book says n starts at 1 4(6) = 4(5)(15)/6 4(6)(6) = 4(5)(15) 4(2.3)(2.3) = 4(5)(5.3) (2)(3)(2) = (5)(5) doesnt seem to work for n=1 does it?

OpenStudy (anonymous):

no it doesn't, oh okay i was supposed to put the one in the two parts equal to each other

OpenStudy (amistre64):

well, for n=1 we only have those parts yes and it turns out that in order for it to actually work, in the basis step, n has to be a complex number :)

OpenStudy (anonymous):

so this statement is false because they don't equal each other?

OpenStudy (amistre64):

correct, there is at least one value of n that this simply does not hold true. And since it is not true for ALL cases, it either needs to be amended or tossed out.

OpenStudy (anonymous):

Okay I have a question, 4(6) = 4(5)(15)/6 4(6)(6) = 4(5)(15) <<<---What did you do to get from here 4(2.3)(2.3) = 4(5)(5.3) <<<---To here? Divide by six?

OpenStudy (amistre64):

multiplied by 6; but that is immaterial. I was simply reworking it to its common factors to show that they do not exist.

OpenStudy (amistre64):

simply letting n =1 to get 24 = 50 is an absurdity enough to claim its false

OpenStudy (anonymous):

okay got it. so for this, 1^2 + 4^2 + 7^2 + ...+ (3n - 2)^2 = n(6n^2 - 3n - 1) I would take the two parts equal to each other and put one in for n?

OpenStudy (amistre64):

n=1 is the basis step, if its is true for n=1, we can then show that it is true for n=2, 3, 4, 5, ... inductively so, establish that (3n - 2)^2 = n(6n^2 - 3n - 1) is true for n=1

OpenStudy (amistre64):

true or false that is; if true, there is more work to do :)

OpenStudy (anonymous):

(3n - 2)^2 = n(6n^2 - 3n - 1)/2 (3(1) - 2)^2 = 1(6(1)^2 - 3(1) - 1)/2 (3 - 2)^2 = 1((6)^2 - 3 - 1)/2 1^2 = 1(36 - 3 - 1)/2 = 1(32)/2 = 32/2 1^2 = 16 Seems like it doesn't work, unless I did something wrong...

OpenStudy (amistre64):

theres a /2? (3n - 2)^2 = n(6n^2 - 3n - 1)/2 (3 - 2)^2 = (6 - 3 - 1)/2 1 = (3 - 1 )/2 1 = 2/2 , then its true for n=1

OpenStudy (amistre64):

yeah 6(1)^2 is not equal to (6)^2 :)

OpenStudy (amistre64):

now to make it more general, let n=k and rewrite the setup, this way we can avoid n=1 for the moment and assume it is true for some n=k. 1^2 + 4^2 + 7^2 + ...+ (3k - 2)^2 = k(6k^2 - 3k - 1)/2 add on the next (k+1) formula to each side ... ...+ (3k - 2)^2 + (3(k+1) - 2)^2 = k(6k^2 - 3k - 1)/2 + (3(k+1) - 2)^2 ^^^^^^^^^^^ ^^^^^^^^^^^ now algebra the right hand side to LOOK like the original setup, but with (k+1): (k+1) (6(k+1)^2 - 3(k+1) - 1)/2 if we can work it back to look like this, then we know that it works for some k, and for some k+1 .... always. we know it works for k=n=1, therefore it has to work for n=2,3,4,5,...

OpenStudy (amistre64):

so .. get to work :) k(6k^2 - 3k - 1)/2 + (3(k+1) - 2)^2, transformed into (k+1) (6(k+1)^2 - 3(k+1) - 1)/2

OpenStudy (anonymous):

I'm sorry, I'm just confused, I got that n=k, then from there, I get...(3k - 2)^2 = k(6k^2 - 3k - 1)/2 Then what do I do with this (k + 1)

OpenStudy (anonymous):

@amistre64

OpenStudy (amistre64):

ill be right back ... got to change computers

OpenStudy (anonymous):

ok

OpenStudy (amistre64):

k+1 is what we need to prove; so we need to add the next "kth" value to both sides of the setup. what you add to one side you add to the other. we know that for each new term on the left, it follows the expression: (3k - 2)^2 so, the (k+1)th term has the same form, but with k+1 instead of k.

OpenStudy (anonymous):

If you could do one line at a time that would be good

OpenStudy (amistre64):

so, to prove it is good for the (k+1)th term; add (3(k+1) - 2)^2 to each side

OpenStudy (amistre64):

gotta move to the library, getting kicked out of my office space :)

OpenStudy (anonymous):

lol okay

OpenStudy (anonymous):

so then I get, (3(k + 1) - 2)^2 + (3k - 2)^2 = k(6k^2 - 3k - 1)/ + (3(k + 1) -2)^2

OpenStudy (amistre64):

you almost see it ..... we already have it setup that: 1^2 + 4^2 + 7^2 + ...+ (3k - 2)^2 = k(6k^2 - 3k - 1)/2 to clear the clutter, lets shove that left hand side into some P(k) P(k) = k(6k^2 - 3k - 1)/2 now, P(k+1) = P(k) + (3(k+1) - 2)^2 or simply, P(k+1) = k(6k^2 - 3k - 1)/2 + (3(k+1) - 2)^2

OpenStudy (amistre64):

all we really can say about this is that the "FORM" is correct: [*] (6[*]^2 - 3[*] - 1) ------------------- 2 we want to show the P(k+1) fits that form

OpenStudy (amistre64):

\[k\frac{(6k^2 - 3k - 1)}{2} + (3(k+1) - 2)^2\] \[\frac{k(6k^2 - 3k - 1) + 2(3(k+1) - 2)^2}{2}\] \[\frac{6k^3 - 3k^2 - k + 2(3k+1)^2}{2}\] \[\frac{6k^3 - 3k^2 - k + 2(9k^2+6k+1)}{2}\] \[\frac{6k^3 - 3k^2 - k + 18k^2+12k+2}{2}\] \[\frac{6k^3 + 15k^2+11k+2}{2}\] we need to factor out a k+1 from the top now , if possible 6k^3 + 15k^2+11k+2 0 -6 -9 -2 ---------------------- -1 | 6 9 2 0 \[\frac{(k+1)(6k^2 + 9k+2)}{2}\] dbl chk my work, im sure i messed up someplace along the way

OpenStudy (anonymous):

in the second line where did you get the 2?

OpenStudy (amistre64):

adding fractions .... 1/2 + a = (1+2a)/2

OpenStudy (amistre64):

it might be helpful to try to get the finished outcome to this stage :) that way we can see a clearer way to get to it

OpenStudy (amistre64):

\[\frac{(k+1)(6(k+1)^2 - 3(k+1) - 1)}{2} \] \[\frac{(k+1)(6(k^2+2k+1) - 3k-3 - 1)}{2} \] \[\frac{(k+1)(6k^2+12k+6 - 3k-4)}{2} \] \[\frac{(k+1)(6k^2+9k+2)}{2} \]

OpenStudy (amistre64):

thats was simpler lol soo, now we have a path to get from that stumper to the form we want ...

OpenStudy (amistre64):

\[\frac{(k+1)(6k^2+9k+2)}{2}\] \[\frac{(k+1)(6k^2+(12-3)k+(6-4))}{2}\] \[\frac{(k+1)(6k^2+12k -3k+6-4)}{2}\] \[\frac{(k+1)(6(k^2+2k +1)-3k-3-1))}{2}\] \[\frac{(k+1)(6(k+1)^2-3(k+1)-1)}{2}\] tada!! lol

OpenStudy (amistre64):

all this does, is show us that if it works for some integer, k; then it also works for the k+1 integer as well. we know it works for k=n=1; so it has to work for 1+1, for 1+1+1, for 1+1+1+1 .... etc

OpenStudy (amistre64):

third line might be better if we swap the -3k and +6 to show that we then complete the square .... had it in my head :)

OpenStudy (amistre64):

i have to go sit in class for 3 hours, so have fun with it ... think over how i did it and work it out in your own mind. good luck

OpenStudy (anonymous):

thanks